# Test-time Correction with Human Feedback: An Online 3D Detection System via Visual Prompting

### v2

@@ Abstract
This paper introduces Test-time Correction (TTC) system, a novel online 3D detection system designated for online correction of test-time errors via human feedback, to guarantee the safety of deployed autonomous driving systems. Unlike well-studied offline 3D detectors frozen at inference, TTC explores the capability of instant online error rectification. By leveraging user feedback with interactive prompts at a frame, e.g., a simple click or draw of boxes, TTC could immediately update the corresponding detection results for future streaming inputs, even though the model is deployed with fixed parameters. This enables autonomous driving systems to adapt to new scenarios immediately and decrease deployment risks reliably without additional expensive training. To achieve such TTC system, we equip existing 3D detectors with Online Adapter (OA) module, a prompt-driven query generator for online correction. At the core of OA module are visual prompts, images of missed object-of-interest for guiding the corresponding detection and subsequent tracking. Those visual prompts, belonging to missed objects through online inference, are maintained by the visual prompt buffer for continuous error correction in subsequent frames. By doing so, TTC consistently detects online missed objects and immediately lowers driving risks. It achieves reliable, versatile, and adaptive driving autonomy. Extensive experiments demonstrate significant gain on instant error rectification over pre-trained 3D detectors, even in challenging scenarios with limited labels, zero-shot detection, and adverse conditions. We hope this work would inspire the community to investigate online rectification systems for autonomous driving post-deployment. Code would be publicly shared.
Figure 1: TTC system could instantly respond to missed objects via human feedback, detecting and tracking these targets in subsequent frames consistently. As depicted on the left, human feedback might come from any views of target objects, including missed objects in the current and/or previous frames, as well as diverse viewpoints across scenes, styles, sources, and poses. Such a system improves offline-trained 3D detectors by rectifying online driving behavior immediately, reducing safety risks via test-time correction. TTC enables a reliable and adaptive online autonomous driving system.
Equal contribution.
@@ 1 Introduction
Visual-based 3D object detection, which localizes and classifies 3D objects from visual imagery, plays a crucial role in autonomous driving systems. Visual autonomous driving frameworks [20, 19, 3, 9] rely heavily on accurate 3D detection outcomes to predict future driving behaviors and plan the trajectory of the ego vehicle. Existing 3D object detectors [27, 62, 35, 46, 21] typically follow an offline training and deployment pipeline. Once the model is trained and deployed on self-driving cars, it is expensive to update new behaviors, e.g., another turn of offline re-training or fine-tuning. That is, when the system fails to perceive important objects or fails in novel scenarios due to a domain shift, these offline solutions cannot update themselves online to rectify mistakes immediately and detect missed objects. Such a caveat poses significant safety risks to reliable driving systems, e.g., dangerous driving behaviors such as improper lane changing, turning, or even collisions.
@@
To guarantee safety, we argue that 3D detectors deployed on autonomous driving systems can rectify missed detection on the fly during test time. As depicted in Figure 2 (left), for error rectification, existing 3D detectors rely on offline pipelines, encompassing a full-suite procedure of data collection, annotation, training, and deployment. This requires significant human workloads and resources for labeling and re-training, days or even weeks to fulfill. Other than the offline pipeline to improve models, we desire deployed 3D detectors also capable of test-time correction since such delays in offline updating are unacceptable when facing risks on the road, where safety is of the utmost priority.
@@
In this work, we explore a new 3D detection system capable of Test-Time error Correction based on human feedback online, namely TTC, akin to how human drivers respond, as shown in Figure 2 (right). The system is designed to address missed detection issues in autonomous driving scenarios, enabling existing 3D detectors to instantly correct errors through online human input, thereby preventing the risks of erroneous decisions caused by missed detections. Inspired by the principles of In-context Learning Customization [58, 43] in large language models (LLMs) [40, 52], we achieve the TTC system by leveraging images of missed objects collected from human feedback as context prompts. These prompts assist deployed 3D detectors in identifying and localizing previously unrecognized objects in later streaming input frames, without extra parameter updates during testing.
@@
The proposed TTC includes two components: Online Adapter (OA) that enables 3D detectors with visual promotable ability, and a visual prompt buffer that records missing objects. The core design is “visual prompts”, the visual object representation derived from human feedback. Existing promptable 3D detection methods typically utilize text, boxes, or clicks as prompts [4, 7, 66, 15, 2]. However, text prompts can be ambiguous and may not describe the target objects effectively. Meanwhile, box and point prompts struggle to handle streaming data. These limitations indicate that such prompts are inadequate for real-time autonomous driving tasks. Visual prompts cover arbitrary imagery views of target objects, i.e., views in different zones, styles, and timestamps (Figure 3), indicating the identity of target objects. With visual prompts, OA module generates corresponding queries, locates corresponding objects within streaming inputs, and facilitates 3D detectors to output 3D boxes.
@@
To enable consecutive error rectification for video streaming, we design a dynamic visual prompt buffer to maintain visual prompts of all past unrecognized objects. In each iteration, we use all visual prompts in this buffer as inputs of the TTC system. This enables the continuous detection and tracking of all previously missed objects, redressing online errors for streaming input effectively. Further, to refrain the buffer from undesirable expansion, we introduce a “dequeue” operation to ensure its bounded size, allowing for consecutive rectification without excessive overhead.
@@
We conduct experiments on the nuScenes dataset [1]. Given the novel setting of our TTC system, the assessment focuses on its abilities for instant online error correction. To this end, we design extensive experiments to verify the key aspects of the overall system: the effectiveness of TTC system in rectifying errors over the online video stream; and the effectiveness of TTC when encountering challenging, extreme scenarios with large amounts of missed detections.
Remarkably, the TTC system significantly improves the offline 3D detectors for test-time performance. Specifically, with test-time rectification, TTC improves offline monocular [67], multi-view [57], and BEV detectors [62] by 5.0%, 12.7%, and 12.1% EDS  [^2] EDS is a class-agnostic version of nuScenes Detection Score (NDS) [1], treating all objects as class-agnostic entities and ignoring the velocity and attribute to evaluate out-of-distribution 3D detection. , respectively. Second, on challenging scenarios, the TTC system exhibits even more substantial gains, with improvements of 14.4%, 21.6%, 13.6%, and 4.7% EDS on tasks like distant 3D detection and vehicle-focused detection with limited annotations, zero-shot extensions, as well as scenarios with domain shifts, respectively.
@@ 2 Related Work
In this paper, we introduce the task of online 3D detection, which focuses on the instant rectification of errors during the online testing phase of 3D detectors. The primary goal is to enable the continuous detection of objects missed by offline-trained 3D detectors, without requiring additional training. This work is closely related to several areas, including 3D detection, tracking, test-time adaption and interactive vision models. Here, we discuss 3D detection and tracking, while leave others in the supplementary material Appendix B.
@@
3D Object Detection. 3D object detection is a cornerstone task in autonomous driving. Traditional 3D detectors primarily focus on the performance of pre-trained models [50, 25, 65, 44, 27]. Once deployed, the model’s output is fixed during inference, with no ability to intervene in real time to correct missed detections. In recent years, some works have explored instruction-based detection systems, offering interfaces that accept human prompts [4, 7, 66, 15, 70, 2, 55]. However, these systems typically rely on boxes, points, or language as interfaces. Boxes and points struggle to handle temporal data [7, 66, 15], while language prompts often depend on LLMs [4, 70, 2, 55], leading to higher latency and potential ambiguities. In contrast, online 3D detection focuses on real-time error correction during sequential detection. By utilizing visual prompts, we enable accurate correction of target objects in streaming data instantly.
@@
Target Object Tracking. Tracking is another direction closely related to online 3D detection [28, 41, 42, 68, 18]. Since our approach uses visual prompts to identify targets, it closely resembles single object tracking (SOT) in terms of its objectives [61, 5, 10, 11]. While current SOT methods based on images are limited to generating 2D detection outputs and presume the object query as the initial frame of the ongoing video, our TTC system stands out. It can directly identify and track the associated 3D bounding boxes using visual prompts, even when these prompts originate from varied sources, scenes, and zones with distinct styles.
3 Test-time Correction with Human Feedback
@@
In this section, we elaborate on our TTC, an online test-time error rectification system for 3D detection to detect and track missed objects during on-road inference with the guidance of human feedback. We start with an overview in Section 3.1, then delve into OA module and the visual prompt buffer in Section 3.2, and Section 3.3, respectively.
@@ 3.1 Overview
We convert online human feedback, i.e., clicks, boxes, or images from the Internet, into a uniform representation called “visual prompts”, image descriptions of target objects. Such image-based descriptions can cover arbitrary views of objects, including pictures taken from diverse zones, weather, timestamps, or even from out-of-domain sources such as stylized Internet images. Upon visual prompts, the TTC system, a recurrent framework, is designed to engage in sustained interaction with human users, continuously learning to detect and track new objects. As Figure 4 (left) shows, it comprises two key components: 1) TTC-3D Detector, any 3D detector equipped with OA module for in-context 3D detection and tracking via visual prompts, and 2) an extendable visual prompt buffer storing visual prompts of all previously missed objects, enabling continuous online error rectification.
During online inference after being deployed on cars, whenever an error occurs, i.e., miss an object, users can add the unrecognized object to visual prompt buffer by clicking on it in the image. The model then detects the corresponding 2D boxes based on the user-provided clicking prompt and updates the visual prompt buffer with the associated image patch. In subsequent frames, TTC-3D Detector leverages the stored visual prompts to detect and track previously missed 3D objects. This enables instant error correction and continuous improvement of 3D detection during online operation.
@@ 3.2 Online Adapter (OA)
OA module is designed as a plug-and-play module. It receives human prompts and transforms them into queries that can be processed by traditional detectors. When trained with OA module, traditional detectors can understand human prompts to produce detection results, enabling online error correction without the need for parameter updates.
Prompt-driven Query Generating. As shown in Figure 4 (right), OA module is flexible to handle four prompts in different forms: object query prompts for traditional offline 3D detection, box and point prompts for collecting test-time feedback, and visual prompts for consistent error correction. Specifically, these prompts are processed as follows:
  • For object prompts $\mathcal{P}_{o}$, the OA module generates a set of learnable embeddings as queries, akin to traditional 3D detectors, which are updated during training as demonstrated in previous works [27, 56, 67].
  • For box $\mathcal{P}_{b}$ and point $\mathcal{P}_{p}$ prompts, the OA module encodes them with their location and shape, representing them as Fourier features [51].
  • For visual prompts $\mathcal{P}_{v}$, the OA module first extracts the visual prompt features $\mathcal{Z}_{v}$ by an encoder [17], then localizes their corresponding objects within the input images $\mathcal{X}_{v}$, and finally adds their features with the Fourier positional encoding as subsequent inputs:
@@
Visual Prompt Alignment. We perform the Align operation to localize target objects in the input images, by visual prompts. To handle flexible visual prompts, which can be image descriptions of objects in any view, scene, style, or timestamp, we employ contrastive mechanisms [59, 16] as the key design to retrieve target objects at different styles. This allows the module to detect the target objects effectively, even when they exhibit diverse visual styles and appearances.
Specifically, we use two multi-layer perceptrons (MLPs) to first align the channels of image features $\mathcal{Z}$ and visual prompt features $\mathcal{Z}_{v}$. We then compute the dot product between the aligned feature maps to obtain a similarity map. To further retrieve the coordinates of target objects from the similarity map, we multiply it with the original image features $\mathcal{Z}$, and apply another MLP with channels of 2 to regress the spatial positions $\mathcal{X}_{v}$ of target objects.
Instance Ambiguity & Loss. Sometimes, visual prompts might exhibit instance ambiguity, where multiple objects in the image match the visual descriptions of prompts. For example, suppose the visual prompt describes a traffic cone and several similar-looking traffic cones present in the input images. It can be challenging to uniquely identify the specific object-of-interest.
@@
For similarity supervision, we generate binary segmentation labels based on the ground-truth 2D boxes. Focal loss [30] and Dice loss [37] are used for optimization. To supervise the visual prompt localization $\mathcal{X}_{v}$, we use the Smooth-$l$1 loss with the target as the center coordinates of ground-truth 2D bounding boxes. For dealing with instance ambiguity, we refer to SAM [24] and only backpropagate the sample with the minimum localization loss during each training iteration. For more details, please refer to the supplementary material Section C.2.
Model Design. The overall mechanism of TTC-3D Detector is depicted in Figure 4 (right). Based on any traditional offline-trained 3D detector, BEV detector, or monocular detector, we integrate OA module and train it to be promptable. Specifically, OA module takes features extracted by the image encoder of the corresponding 3D detector, along with various forms of prompts as inputs. It encodes these prompts and generates a series of queries, represented as $\mathcal{P}=\{\mathcal{P}_{o};<\mathcal{P}_{b},\mathcal{P}_{p},\mathcal{P}_{v}>\}$, where <>expectation<...>< … > denotes an arbitrary combination of different prompts. These queries are then fed into the transformer decoder of the 3D detector to output 3D boxes following human feedback.

### v3

# Test-time Correction: An Online 3D Detection System via Visual Prompting

@@ Abstract
This paper introduces Test-time Correction (TTC), an online 3D detection system designed to rectify test-time errors using various auxiliary feedback, aiming to enhance the safety of deployed autonomous driving systems. Unlike conventional offline 3D detectors that remain fixed during inference, TTC enables immediate online error correction without retraining, allowing autonomous vehicles to adapt to new scenarios and reduce deployment risks. To achieve this, we equip existing 3D detectors with an Online Adapter (OA) module—a prompt-driven query generator for real-time correction. At the core of OA module are visual prompts: image-based descriptions of objects of interest derived from auxiliary feedback such as mismatches with 2D detections, road descriptions, or user clicks. These visual prompts, collected from risky objects during inference, are maintained in a visual prompt buffer to enable continuous correction in future frames. By leveraging this mechanism, TTC consistently detects risky objects, achieving reliable, adaptive, and versatile driving autonomy. Extensive experiments show that TTC significantly improves instant error rectification over frozen 3D detectors, even under limited labels, zero-shot settings, and adverse conditions. We hope this work inspires future research on post-deployment online rectification systems for autonomous driving.
Index Terms:
3D Object Detection, Test-time correction, Online detection system.
@@ 1 Introduction
Visual-based 3D object detection, which localizes and classifies 3D objects from visual imagery, plays a crucial role in autonomous driving systems. Visual autonomous driving frameworks [hu2023_uniad, hu2022stp3, casas2021mp3, Cui_2021] rely heavily on accurate 3D detection outcomes to predict future driving behaviors and plan the trajectory of the ego vehicle. Existing 3D object detectors [li2022bevformer, yang2023bevformerV2, liu2022petr, CaDDN, huang2021bevdet] typically follow an offline training and deployment pipeline. Once the model is trained and deployed on self-driving cars, it is expensive to update new behaviors, e.g., another turn of offline re-training or fine-tuning. That is, when the system fails to perceive important objects or fails in novel scenarios due to a domain shift, these offline solutions cannot update themselves online to rectify mistakes immediately and detect missed objects. Such a caveat poses significant safety risks to reliable driving systems, e.g., dangerous driving behaviors such as improper lane changing, turning, or even collisions.
@@
To guarantee safety, we argue that 3D detectors deployed on autonomous driving systems can rectify missed detection on the fly during test time. As depicted in Figure˜1 (left), for error rectification, existing 3D detectors rely on offline pipelines, encompassing a full-suite procedure of data collection, annotation, training, and deployment. This requires significant human workloads and resources for labeling and re-training, days or even weeks to fulfill. Other than the offline pipeline to improve models, we desire deployed 3D detectors also capable of test-time correction since such delays in offline updating are unacceptable when facing risks on the road, where safety is of the utmost priority.
@@
In this work, we propose a novel 3D detection system, TTC, that provides online Test-Time Correction by leveraging various auxiliary information (e.g., 2D detections, road descriptions, or human feedback), as shown in Figure˜1 (right). The system is designed to address missed detection issues in autonomous driving scenarios, enabling existing 3D detectors to leverage various real-time feedback from multiple sources for instant error correction, thereby reducing the likelihood of erroneous decisions caused by undetected objects. Inspired by the principles of In-context Learning Customization [cot, llm2023peng] in large language models (LLMs) [gpt4, gemini], we implement test-time correction by converting auxiliary information into prompts and feeding them into the system for online adaptation.
@@
The proposed TTC includes two components: Online Adapter (OA) that enables 3D detectors with visual promotable ability, and a visual prompt buffer that records risky objects. The core design is “visual prompts”, the visual object representation derived from various auxiliary sources. Existing promptable 3D detection methods typically utilize text, boxes, or clicks as prompts [chen2024ll3da, choi2024idet3d, zhang2023sam3d, guo2024sam2point, cao2024coda]. However, text prompts often require integration with large language models, and supplying a text prompt on a per-frame basis results in significant latency. Meanwhile, box and point prompts on specific frames struggle to handle streaming data. These limitations indicate that such prompts are inadequate for real-time autonomous driving tasks. Visual prompts cover arbitrary imagery views of target objects, i.e., views in different zones, styles, and timestamps (Figure˜2), indicating the identity of target objects. Once generated, visual prompts can operate on subsequent frames, accurately describing and tracking target risky objects. With visual prompts, OA module generates corresponding queries, locates corresponding objects within streaming inputs, and facilitates 3D detectors to output 3D boxes.
@@
To enable consecutive error rectification for video streaming, we design a dynamic visual prompt buffer to maintain visual prompts of all risky objects. In each iteration, we use all visual prompts in this buffer as inputs to the TTC system. This enables the continuous detection and tracking of all risky objects, redressing online errors for streaming input effectively. Further, to prevent the buffer from undesirable expansion, we introduce a “dequeue” operation to ensure its bounded size, allowing for consecutive rectification without excessive overhead.
@@
We conduct experiments on the nuScenes dataset [nuscenes2019]. Given the novel setting of our TTC system, the assessment focuses on its abilities for instant online error correction. To this end, we design extensive experiments to verify the key aspects of the overall system: the effectiveness of TTC system in rectifying errors over the online video stream; and the effectiveness of TTC when encountering challenging, extreme scenarios with large amounts of missed detections.
Remarkably, the TTC system significantly improves the offline 3D detectors for test-time performance. Specifically, with test-time rectification, TTC improves offline monocular [zhang2022monodetr], multi-view [wang2023object], and BEV detectors [yang2023bevformerV2] by 2.4%, 9.7%, and 9.3% mAP, respectively. Second, on challenging scenarios, the TTC system exhibits even more substantial gains, with improvements of 10.1%, 20.0%, 7.2%, and 3.8% mAP on tasks like distant 3D detection and vehicle-focused detection with limited annotations, zero-shot extensions, as well as scenarios with domain shifts, respectively.
@@ 2 Related Work
In this paper, we introduce the task of online 3D detection, which focuses on the instant rectification of errors during the online testing phase of 3D detectors. The primary goal is to enable the continuous detection of objects missed by offline-trained 3D detectors, without requiring additional training after deployment. This work is closely related to several areas, including 3D detection, tracking, test-time adaptation, and interactive vision models.
2.1 3D Object Detection
@@
3D object detection is a cornerstone task in autonomous driving. Traditional 3D detectors primarily focus on the performance of pre-trained models [PVRCNN, lang2018pointpillars, yin2021center, philion2020lift, li2022bevformer]. Once deployed, the model’s output is fixed during inference, with no ability to intervene in real time to correct missed detections. In recent years, some works have explored instruction-based detection systems, offering interfaces that accept prompts [chen2024ll3da, choi2024idet3d, zhang2023sam3d, guo2024sam2point, zhu2025unifying, cao2024coda, wang2024omnidrive]. However, these systems typically rely on boxes, points, or language as interfaces. Boxes and points struggle to handle temporal data [choi2024idet3d, zhang2023sam3d, guo2024sam2point], while language prompts often depend on LLMs [chen2024ll3da, zhu2025unifying, cao2024coda, wang2024omnidrive], leading to higher latency and potential ambiguities. In contrast, online 3D detection focuses on real-time error correction during sequential detection. By utilizing visual prompts, we enable accurate correction of target objects in streaming data instantly.
2.2 Target Object Tracking
@@
Tracking is another direction closely related to online 3D detection [pnpnet, pang2021simpletrack, Multiple3Dtracking, zhang2022mutr3d, Hu2021QD3DT]. Since our approach uses visual prompts to identify targets, it closely resembles single object tracking (SOT) in terms of its objectives [yan2021learning, chen2023seqtrack, cui2022mixformer, cui2024mixformerv2]. While current SOT methods based on images are limited to generating 2D detection outputs and presume the object query as the initial frame of the ongoing video, our TTC system stands out. It can directly identify and track the associated 3D bounding boxes using visual prompts, even when these prompts originate from varied sources, scenes, and zones with distinct styles.
2.3 Interactive Vision Models
Interactive vision models are designed for tasks based on user inputs. As one of the fundamental tasks in computer vision, they are extensively researched with numerous breakthroughs [lazysnap, focalclick, simpleclick, deepinteract, randomwalk]. Particularly, the advent of the Segment Anything Model (SAM) [kirillov2023segany] has sparked a surge of progress, with applications spanning reconstruction [shen2023anything3d], detection [ren2024grounded, yang2023track], segmentation [zou2023segment], image editing [gao2023editanything], and more. Compared to existing models, which typically rely on prompts such as clicks, boxes, or scribbles, in this work, we study visual prompts, a new prompt referring to the actual images of objects with arbitrary poses and styles. Visual prompts enable continuous detection and tracking of target objects in streaming videos, facilitating the immediate correction of failed detections at test time. This represents a departure from traditional prompt types, offering more natural and dynamic interactions with the visual content.
2.4 Test-time Adaptation
Test-time adaptation aims to improve model performance on test data by adapting the model to test samples, even in the presence of data shifts [zhang2023adanpc, niu2022efficient, niutowards, schneider2020improving, wang2020tent, liu2021ttt++]. Most TTA methods update network parameters during testing—via entropy minimization, self-supervised objectives, or batch-norm statistic adaptation. More recently, several works have brought TTA to object detection [ruan2024fully, vs2023towards, cao2024exploring], and to 3D detection in particular (e.g., MonoTTA [lin2025monotta], Reg-TTA3D [yuan2024reg], and MOS [chenmos]). These approaches typically fine-tune on streaming target batches across domains or perform input-conditioned model selection/mixture, still relying on weight updates or model switching at test time. In contrast, our method does not modify the original model parameters. Instead, it performs parameter-free, prediction-level rectification driven by various auxiliary feedback, enabling instance-level online correction without retraining or compromising source-domain accuracy.
3 Test-time Correction
@@
In this section, we elaborate on our TTC, an online test-time error rectification system for 3D detection to detect and track missed objects during on-road inference with the guidance of various feedback. We start with an overview in Section˜3.1, then delve into OA module and the visual prompt buffer in Section˜3.2, and Section˜3.3, respectively.
@@ 3.1 Overview
We convert various online auxiliary feedback, i.e., such as human-provided clicks and boxes, road descriptions, and misaligned detection results compared with 2D detectors, into a uniform representation called “visual prompts”, image descriptions of missed or risky objects. Such image-based descriptions can cover arbitrary views of objects, including pictures taken from diverse zones, weather, timestamps, or even from out-of-domain sources such as stylized Internet images. Based on visual prompts, the TTC system is designed as a real-time framework capable of continuously detecting and tracking risky objects, even when the raw feedback may suffer from high latency. As Figure˜3 (left) shows, it comprises two key components: (i) TTC-3D Detector, any 3D detector equipped with OA module for in-context 3D detection and tracking via visual prompts, and (ii) an extendable visual prompt buffer storing visual prompts of all previously missed objects, enabling continuous online error rectification.
During online inference after deployment in vehicles, the model can leverage auxiliary information by converting it into visual prompts and adding them to the visual prompt buffer. For user clicks, the model first detects the corresponding 2D boxes based on the user-provided point prompt and updates the visual prompt buffer with the associated image patch. For user-drawn boxes or cases involving misaligned 2D detection results, the corresponding image regions can be directly pre-processed via cropping. For road descriptions, relevant visual prompts can be retrieved via online search based on the provided textual descriptions. TTC-3D Detector leverages the stored visual prompts to detect and track risky 3D objects. This enables instant error correction and continuous improvement of 3D detection during online operation.
@@ 3.2 Online Adapter (OA)
OA module is designed as a plug-and-play module. It receives prompts and transforms them into queries that can be processed by traditional detectors. When trained with OA module, traditional detectors can understand various prompts to produce detection results, enabling online error correction without the need for parameter updates.
Prompt-driven Query Generating. As shown in Figure˜3 (right), OA module is flexible to handle four prompts in different forms: object query prompts for traditional offline 3D detection, box and point prompts for feedback converting, and visual prompts for consistent error correction. Specifically, these prompts are processed as follows:
  • For object prompts $\mathcal{P}_{o}$, the OA module generates a set of learnable embeddings as queries, akin to traditional 3D detectors, which are updated during training as demonstrated in previous works [li2022bevformer, detr3d, zhang2022monodetr].
  • For box $\mathcal{P}_{b}$ and point $\mathcal{P}_{p}$ prompts, the OA module encodes them with their location and shape, representing them as Fourier features [tancik2020fourfeat].
  • For visual prompts $\mathcal{P}_{v}$, the OA module first extracts the visual prompt features $\mathcal{Z}_{v}$ by an encoder [ResNet], then localizes their corresponding objects within the input images $\mathcal{X}_{v}$, and finally adds their features with the Fourier positional encoding as subsequent inputs:
@@
Visual Prompt Alignment. We perform the Align operation to localize target objects in the input images, by visual prompts. To handle flexible visual prompts, which can be image descriptions of objects in any view, scene, style, or timestamp, we employ contrastive mechanisms [wu2018unsupervised, he2020moco] as the key design to retrieve target objects at different styles. This allows the module to detect the target objects effectively, even when they exhibit diverse visual styles and appearances.
Figure 4: Concrete mechanism of visual prompt alignment. This figure illustrates monocular input. When multi-view images are employed, this alignment operation flattens the different views and still generates NN peak candidate positions.
As described in Figure˜4, given the visual prompt $\mathcal{P}_{v}$, we first extract the visual prompt features $\mathcal{Z}_{v}$ using a lightweight encoder, e.g., ResNet18 [ResNet] in our implementation. Then, we use two separate 2-layer perceptrons, each with 128 and 64 output channels, to align the channel dimensions of the prompt features and the input image features. This results in the aligned feature maps $\hat{\mathcal{Z}_{v}}\in\mathbb{R}^{M\times 64}$ and $\hat{\mathcal{Z}}\in\mathbb{R}^{KHW\times 64}$, where $M$ is the number of visual prompts, $K$ is the number of image views, $H$ and $W$ are the spatial dimensions of image features. Finally, we compute the cosine similarity ($\odot$ in Figure˜4) between $\hat{\mathcal{Z}}_{v}$ and $\hat{\mathcal{Z}}$ to obtain the similarity map $\mathcal{S}\in\mathbb{R}^{M\times KHW}$, which encodes the alignments between the visual prompts and image features.
Instance Ambiguity & Loss. Sometimes, visual prompts might exhibit instance ambiguity, where multiple objects in the image match the visual descriptions of prompts. For example, suppose the visual prompt describes a traffic cone and several similar-looking traffic cones present in the input images (highlighted by “orange box” in Figure˜4). It can be challenging to uniquely identify the specific object of interest.
@@
For similarity supervision, we generate binary segmentation labels based on the ground-truth 2D boxes. Focal loss [FocalLoss] and Dice loss [Milletari2016vnet] are used for optimization. To supervise the visual prompt localization $\mathcal{X}_{v}$, we use the Smooth-ll1 loss with the target as the center coordinates of ground-truth 2D bounding boxes. For dealing with instance ambiguity, we refer to SAM [kirillov2023segany] and only backpropagate the sample with the minimum localization loss during each training iteration.
Model Design. The overall mechanism of TTC-3D Detector is depicted in Figure˜3 (right). Based on any traditional offline-trained 3D detector, BEV detector, or monocular detector, we integrate OA module and train it to be promptable. Specifically, OA module takes features extracted by the image encoder of the corresponding 3D detector, along with various forms of prompts as inputs. It encodes these prompts and generates a series of queries, represented as $\mathcal{P}=\{\mathcal{P}_{o};<\mathcal{P}_{b},\mathcal{P}_{p},\mathcal{P}_{v}>\}$, where <><...> denotes an arbitrary combination of different prompts. These queries are then fed into the transformer decoder of the 3D detector to output 3D boxes.